# VideoGuide: Improving Video Diffusion Models without Training Through a Teacher’s Guide

### v2

@@ Appendix C User Study
  • Question: Do you think the videos reflect the given text condition well?
    (5: Strongly Agree / 4: Agree / 3: Neutral / 2: Disagree / 1: Strongly Disagree)
@@
  • Question: Do you think the video’s overall quality is good? (rich detail, unchanging objects)
    (5: Strongly Agree / 4: Agree / 3: Neutral / 2: Disagree / 1: Strongly Disagree)
@@
  • Question: Do you think the video’s overall motion is smooth and dynamic?
    (5: Strongly Agree / 4: Agree / 3: Neutral / 2: Disagree / 1: Strongly Disagree)
Method Text Alignment Overall Quality Smooth And Dynamic Motion
@@
Table 5: User Study. Bold: best, underline: second best.
Tab. 5 shows that our method surpasses the baseline and previous work in all evaluated aspects.
Appendix D Pseudo Code
Pseudo codes regarding our algorithm are provided in the following page.
Algorithm 1 VideoGuide with Sampling Diffusion Model
1:  Initialize ${\boldsymbol{z}}_{T}\sim\mathcal{N}(0,{\mathbf{I}})$
2:  for $t=T,\dots,1$ do
3:     $\hat{{\boldsymbol{\epsilon}}}_{\theta}({\boldsymbol{z}}_{t},t)={\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{z}}_{t},t,\phi)+\lambda[{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{z}}_{t},t,c)-{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{z}}_{t},t,\phi)]$
4:     ${\boldsymbol{z}}_{0|t}=({\boldsymbol{z}}_{t}-\sqrt{1-\bar{\alpha}_{t}}\hat{{\boldsymbol{\epsilon}}_{\theta}}({\boldsymbol{z}}_{t},t))/\sqrt{\bar{\alpha}_{t}}$
5:     ${\boldsymbol{z}}_{t}=\sqrt{\bar{\alpha}_{t}}{\boldsymbol{z}}_{0|t}+\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}},\quad\textit{where}\quad{\boldsymbol{\epsilon}}\sim N(0,{\mathbf{I}})$
6:     if $T-t<I$ then
7:        for $j=0,\dots,\tau$ do
8:           ${\boldsymbol{z}}_{t-j-1}=\sqrt{\bar{\alpha}_{t-j-1}}{\boldsymbol{z}}_{0|t-j}+\sqrt{1-\bar{\alpha}_{t-j-1}}{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{z}}_{t-j},t-j,\phi)$
9:        end for
10:        ${\boldsymbol{z}}^{\prime}_{0|t}=\beta\cdot{\boldsymbol{z}}_{0|t}+(1-\beta)\cdot{\boldsymbol{z}}_{0|t-\tau}$
11:        ${\boldsymbol{z}}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}{\boldsymbol{z}}^{\prime}_{0|t}+\sqrt{1-\bar{\alpha}_{t-1}}{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{z}}_{t},t,\phi)$
12:        ${\boldsymbol{z}}_{t-1}=\textit{LPF}_{\gamma}({\boldsymbol{z}}_{t-1})+\textit{HPF}_{\gamma}({\boldsymbol{\epsilon}}),\quad\textit{where}\quad{\boldsymbol{\epsilon}}\sim N(0,{\mathbf{I}})$
13:     else
14:        ${\boldsymbol{z}}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}{\boldsymbol{z}}_{0|t}+\sqrt{1-\bar{\alpha}_{t-1}}{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{z}}_{t},t,\phi)$
15:     end if
16:  end for
17:  Output: Final video ${\boldsymbol{z}}_{0}$
Algorithm 2 VideoGuide with Guiding Diffusion Model
1:  Initialize ${\boldsymbol{z}}_{T}\sim\mathcal{N}(0,{\mathbf{I}})$
2:  for $t=T,\dots,1$ do
3:     $\hat{{\boldsymbol{\epsilon}}}_{\theta}({\boldsymbol{z}}_{t},t)={\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{z}}_{t},t,\phi)+\lambda[{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{z}}_{t},t,c)-{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{z}}_{t},t,\phi)]$
4:     ${\boldsymbol{z}}_{0|t}=({\boldsymbol{z}}_{t}-\sqrt{1-\bar{\alpha}_{t}}\hat{{\boldsymbol{\epsilon}}_{\theta}}({\boldsymbol{z}}_{t},t))/\sqrt{\bar{\alpha}_{t}}$
5:     ${\boldsymbol{z}}^{(G)}_{t}=\sqrt{\bar{\alpha}^{(G)}_{t}}z_{0|t}+\sqrt{1-\bar{\alpha}^{(G)}_{t}}{\boldsymbol{\epsilon}},\quad\textit{where}\quad{\boldsymbol{\epsilon}}\sim N(0,{\mathbf{I}})$
6:     if $T-t<I$ then
7:        for $j=0,\dots,\tau$ do
8:           ${\boldsymbol{z}}_{0|t-j}^{(G)}=({\boldsymbol{z}}_{t-j}^{(G)}-\sqrt{1-\bar{\alpha}_{t-j}^{(G)}}\hat{{\boldsymbol{\epsilon}}}_{\psi}({\boldsymbol{z}}_{t-j}^{(G)},t-j)/\sqrt{\bar{\alpha}_{t-j}^{(G)}}$
9:           ${\boldsymbol{z}}_{t-j-1}^{(G)}=\sqrt{\bar{\alpha}_{t-j-1}^{(G)}}{\boldsymbol{z}}_{0|t-j}^{(G)}+\sqrt{1-\bar{\alpha}_{t-j-1}^{(G)}}{\boldsymbol{\epsilon}}_{\psi}({\boldsymbol{z}}_{t-j}^{(G)},t-j,\phi)$
10:        end for
11:        ${\boldsymbol{z}}^{\prime}_{0|t}=\beta\cdot{\boldsymbol{z}}_{0|t}+(1-\beta)\cdot{\boldsymbol{z}}_{0|t-\tau}^{(G)}$
12:        ${\boldsymbol{z}}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}{\boldsymbol{z}}^{\prime}_{0|t}+\sqrt{1-\bar{\alpha}_{t-1}}{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{z}}_{t},t,\phi)$
13:        ${\boldsymbol{z}}_{t-1}=\textit{LPF}_{\gamma}({\boldsymbol{z}}_{t-1})+\textit{HPF}_{\gamma}({\boldsymbol{\epsilon}}),\quad where\quad{\boldsymbol{\epsilon}}\sim N(0,{\mathbf{I}})$
14:     else
15:        ${\boldsymbol{z}}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}{\boldsymbol{z}}_{0|t}+\sqrt{1-\bar{\alpha}_{t-1}}{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{z}}_{t},t,\phi)$
16:     end if
17:  end for
18:  Output: Final video ${\boldsymbol{z}}_{0}$
Appendix E More Qualitative Examples

### v3

@@ Appendix E User Study
  • Question: Do you think the videos reflect the given text condition well? (5: Strongly Agree / 4: Agree / 3: Neutral / 2: Disagree / 1: Strongly Disagree)
@@
  • Question: Do you think the video’s overall quality is good? (rich detail, unchanging objects) (5: Strongly Agree / 4: Agree / 3: Neutral / 2: Disagree / 1: Strongly Disagree)
@@
  • Question: Do you think the video’s overall motion is smooth and dynamic? (5: Strongly Agree / 4: Agree / 3: Neutral / 2: Disagree / 1: Strongly Disagree)
Method TA OQ SDM
@@
Table 5: User Study. Text Alignment (TA), Overall Quality (OQ), Smooth and Dynamic Motion (SDM) are evaluated among methods. Bold: best, underline: second best.
Appendix F Application on DiT-based Models
We further evaluate the robustness of our methodology by applying it to different architectures and schedulers. Specifically, we present further evaluation on models that use Diffusion Transformer (DiT) [Peebles2022DiT] architecture: Open-Sora v1.0 [opensora] and Open-Sora v1.2 [opensora]. Each model employs a standard DDIM scheduler (50 steps) and a rectified flow [liu2022flow] scheduler, respectively. In the rectified flow-based configuration, the objective for training can be formulated as follows:
${\bm{z}}_{t}$ $=(1-t){\bm{z}}_{0}+t{\bm{\epsilon}}\quad where\quad t\in[0,1]$ (17)
[…]
Using the objective above we can redefine our method as below:
${\bm{z}}_{0|t_{i}}$ $={\bm{z}}_{t_{i}}+t_{i}\cdot v_{\theta}({\bm{z}}_{t_{i}},t_{i})$ (18)
${\bm{\epsilon}}_{\theta}({\bm{z}}_{t_{i}},t_{i})$ $={\bm{z}}_{t_{i}}-(1-t_{i})\cdot v_{\theta}({\bm{z}}_{t_{i}},t_{i})$
${\bm{z}}_{t_{i-1}}$ $=(1-t_{i-1})f({\bm{z}}_{0|t_{i}},\beta,\tau)+t_{i-1}{\bm{\epsilon}}_{\theta}({\bm{z}}_{t_{i}},t_{i})$
where $f({\bm{z}}_{0|t_{i}},\beta,\tau)$ is the interpolation function between ${\bm{z}}_{0|t_{i}}$ and ${\bm{z}}_{0|t_{i-\tau}}$ with scale $\beta$. The results in Tab. 6 demonstrate that our method improves temporal consistency for both baselines while preserving imaging quality and introducing only a minimal decrease in dynamic degree. These findings indicate that our methodology enhances performance regardless of the underlying architecture and scheduler.
Method Subject consistency ($\uparrow$) Background Consistency ($\uparrow$) Imaging Quality ($\uparrow$) Motion Smoothness ($\uparrow$) Dynamic Degree ($\uparrow$)
OpenSora v1.0 [opensora] (DDIM [song2020denoising]) $0.9735$ $0.9689$ $0.6615$ $0.9678$ $4.97$
OpenSora v1.0 + VideoGuide (self-guided) $0.9763$ $0.9689$ $0.6738$ $0.9754$ $3.88$
OpenSora v1.2 [opensora] (Rectified Flow [liu2022flow]) $0.9725$ $0.9696$ $0.6582$ $0.9881$ $12.68$
OpenSora v1.2 + VideoGuide (self-guided) $0.9808$ $0.9748$ $0.6689$ $0.9903$ $11.07$
Table 6: Quantitative comparison of video generation in DiT-based architecture.
Figure 7: Qualitative Results of VideoGuide on Open-Sora v1.0.
Figure 8: Qualitative Results of VideoGuide on Open-Sora v1.2.
Appendix G Comparison with Orthogonal Methods
A recent study, UniCtrl [chen2024unictrl], attempts to improve semantic consistency and motion quality in an approach orthogonal to ours. In this section, we compare the performance of each technique and assess the feasibility of combining them. Following the recommendation in the paper, we use a motion injection degree of $c=0.2$, while maintaining the same experimental configuration as described in Section D. As illustrated in Table 7, UniCtrl [chen2024unictrl] improves temporal consistency but at the cost of a significant reduction in dynamic degree and imaging quality.
Method Subject consistency ($\uparrow$) Background Consistency ($\uparrow$) Imaging Quality ($\uparrow$) Motion Smoothness ($\uparrow$) Dynamic Degree ($\uparrow$)
AnimateDiff [guo2023animatediff] $0.9183$ $0.9437$ $0.6647$ $0.9547$ $26.67$
AnimateDiff + UniCtrl [chen2024unictrl] $0.9259$ $0.9413$ $0.6032$ $0.9584$ $14.96$
AnimateDiff + Ours $0.9614$ $0.9664$ $0.6671$ $0.9772$ $16.78$
AnimateDiff + UniCtrl + Ours $0.9639$ $0.9628$ $0.5883$ $0.9776$ $5.02$
Table 7: Quantitative comparison with orthogonal methods.
Figure 9: Qualitative Comparison of UniCtrl and VideoGuide.
Appendix H Limitations
While our approach significantly improves the performance of baseline models, it relies on sharing the same Variational Auto-Encoder(VAE) [kingma2013auto] space. In practice, many latent diffusion models utilize the same VAE, making this requirement generally feasible. However, if the VAE spaces differ, one potential solution is to decode, interpolate, and re-encode the features. This process, however, incurs additional computational overhead and risks losing fine details due to iterative encoding-decoding. Developing an effective method to address compatibility across different VAE spaces remains an avenue for future research.
Appendix I More Qualitative Examples
@@ I.1 Prior Distillation
I.2 More Qualitative Comparison Results
Figure 11: More Qualitative Comparison Results of VideoGuide. Top: AnimateDiff with ToonYou, Bottom: AnimateDiff with RCNZCartoon
I.3 More Qualitative Results